OpenStudy (anonymous):

y=x+3, x=11 what does y=

OpenStudy (hoblos):

replace x by 11 in the equation (y=x+3) you get y= 11+3 y=14

OpenStudy (anonymous):

sub x=11 into the equation. y=x+3 y=11+3 y=14

OpenStudy (anonymous):

thank you

OpenStudy (hoblos):

glad to help :))

OpenStudy (anonymous):

.